OpenStudy (anonymous):

My son is taking Trig and it's been a very longtime since I did this can you refresh me on these questions so I can help him understand? 1. Verify each Identity cosx/1-sin(squared)x=sec x 2. sin(squared)xcosxsecx=1-cos(squared)x 3. suppose that sin a= 3/5 and sin B=24/5, where 0 < a < pi/2

OpenStudy (anonymous):

taco wiz khalifa black n yellow

Directrix (directrix):

@pespa You have violated the OpenStudy Code of Conduct •Posting a reply that is not relevant to the topic of discussion within that question.

OpenStudy (whpalmer4):

\[\frac{\cos x}{1-\sin^2 x} = \sec x\]Remember the unit circle: the radius is 1, and cos x is the x distance and sin x is the y distance. Therefore, \(\cos^2x + \sin^2x =1\)Can you use that to your advantage? Also, do you know the definition of \(\sec x\) in terms of the more basic trig functions?

OpenStudy (whpalmer4):

\[\sin^2x\cos x\sec x = 1-\cos^2x\] Replace \(\sec x\) with its basic definition, and also make use of the identity I gave you above.

OpenStudy (anonymous):

ok how are the rest worked?

OpenStudy (whpalmer4):

Sounds like you need to find a good listing of trig identities. There's one I like at http://tutorial.math.lamar.edu/pdf/Trig_Cheat_Sheet.pdf You'll use the identities for the sin and cos of a sum: \[\sin(\alpha\pm\beta) = \sin\alpha\cos\beta\pm\cos\alpha\sin\beta\] \[\cos(\alpha\pm\beta)=\cos\alpha\cos\beta\mp\sin\alpha\sin\beta\]